OpenStudy (anonymous):

Write an equation in point-slope form of the line with slope –8 that contains the point J(–3, 2).

OpenStudy (anonymous):

theres ur answer in there

OpenStudy (anonymous):

do u see it bam loverstar? y-2=-8(x+3) is ur answer

OpenStudy (helder_edwin):

the formula of the slope is \[\large m=\frac{y_1-y_2}{x_1-x_2} \] u have \(m=-8\) and \(J(-3,2)=(x_1,y_1)\). Just replace these two into the equation and u will have the equation.

OpenStudy (anonymous):

yeaa thnks they both helped

OpenStudy (anonymous):

y-2=-8(x+3) is ur answer